OpenStudy (anonymous):

if (x)=12 x+7, what is f^1 (x)

OpenStudy (anonymous):

check is it's f^1(x) or f^-1(x)?

OpenStudy (anonymous):

Are you looking for a derivative? f^1(x), by exponent rule, will be f^1(x) = 12

OpenStudy (anonymous):

this says multiply by 12, add 7 inverse say subtract 7, divide by 12 it is \[f^{-1}(x)=\frac{x-7}{12}\]

OpenStudy (anonymous):

if you like you can write \[y=12x+7\] switch x and y which is what the inverse does and write \[x=12y+7\] solve for y get \[x-7=12y\] \[\frac{x-7}{12}=y\] so \[f^{-1}(x)=\frac{x-7}{12}\] that way